CTD Processing Summary CalCOFI 1101NH CTD Final Data (reformatted 10/2017 for QC columns)
Download 1101 CTD raw cast files zipped Download 1101 FinalQC CTD + bottle data
General CTD Notes - data acquisition cast notes, logistics, processing notes are listed below
CTD sensor corrections derived by comparing 4 secs of CTD sensor data (prior to bottle closure) to bottle samples
Dual T, S, & O2 Primary Sensor Secondary Sensor
Salinity offset (bottle - CTD salinity; > 350m only; Seabird SBE4) -0.0016 -0.0002
Oxygen (ml/L; dual Seabird SBE43) y = 1.0578x - 0.0114
R² = 0.9997
y = 1.0623x - 0.0362  (first 50 stations only; sensor failed on 051-106)
R² = 0.9996
Oxygen (umol/Kg; dual Seabird SBE43) y = 1.049x - 0.4238
R² = 0.9997
y = 1.0538x - 1.4863  (first 50 stations only; sensor failed on 051-106)
R² = 0.9996
Single sensors
Nitrate - ISUS 4sec ave voltage vs Bottle Nitrate (Satlantic MBARI-ISUS v2) y = 29.064x - 2.9761
R² = 0.9903
Please note the ISUS was removed after cast 086 so only casts 001-086 have ISUS estimated nitrate data.
Fluorometer - linear & polynomial regressions

y = 13.439x - 0.2944

R² = 0.7256

y = 4.5466x2 + 11.036x  -0.1847

R² = 0.7329

Primary CTD O2 Sensor vs Bottle O2
Secondary CTD O2 Sensor vs Bottle O2
Primary CTD O2 Sensor vs Bottle O2 Secondary CTD O2 Sensor vs Bottle O2
MBARI-ISUS Voltage vs Bottle NO3 Seapoint Fluorometer Voltage vs Chlorophyll-a
General notes: 
CalCOFI 1101NH: 12 Jan - 06 Feb 2011 on RV New Horizon, 106 stations occupied
Please refer to this webpage for more cruise info:

This cruise was re-processing in June 2015 to add oxygen in umol/Kg for the CTD
database. In the process, JRW noticed the Windows Filter was mis-applied originally
with the ISUS 500 cosine filter applied to voltage 4 instead of 7. This was fixed.
ISUS was only deployed on stations 01-86, removed for 87-106. Also there was a cable
failure on casts 20, 21, 22, 23 so there is no ISUS data on those stations. There
were so ISUS drop-outs during casts 12, 15, 42, 44, 49, & 77. The bad voltages
were replaced by -99 so they would not be used by the station-correction regression
This cruise used archive IEH bottle data to correct CTD sensor data.
It appears that cast 104, the CTD operator did not start data archiving till the
CTD downcast hit ~280m. Refer to the upcast data for surface to 280m data.
More cast notes may be viewed online at:
JRW was not on this cruise.
Cast 001 - 086 had ISUS sensor data
Cast 001 - 083 had nutrient data; plots after this will require cruise-corrected only data
Cast 087 - 106 no ISUS data
Cast 084 - 106 not nutrient data

Cast specific Notes from Console Ops:
Cast 001 - SCCOOS sta 93.4 26.4 at 1630PST due to acoustics calibration; no prodo 1st day.
    Cruise originally mislabeled 1011, fixed in files.
Cast 006 - restarted after naming data file "deck test"
Cast 009 - winch operator tube locked the CTD; 2 bottle lanyards broken & fixed;
    conductive wire looked okay. MGS added yellow tape about cable grip for visibility.
    Wire tension reached ~2000lbs, everything deemed okay after lanyard replacement.
Cast 010 - Oozeki trawl onto station, cod end lost, tow not repeated
Cast 011 - bottles 8 & 11 valves open
Cast 012 - ISUS outage @~170m back on @230m; CTD file edited during 2015 re-process
Cast 018 - two extra bottles closed at surface for O2 test cases
Cast 019 - three extra bottles fired at surface for Megan/LTER
Casts 20, 21, 22, 23 - No ISUS data collected, bad cable but not discovered till
    cast 22 or repaired till cast 24. MGS replaced the ISUS data cable.
Cast 21, 22, 23, & 24 - pH data was good on cast 20 but looked bad on    casts 21, 22, 23 & 24.
Cast 024 - ISUS working, pH not; ship being pushed by strong current, affection sta position
Cast 028 - bottom getting shallow - only 290m to begin then 280...250; CTD depth to 253m
Cast 030 - brown surface water
Cast 031 - was originally not labeled SCCOOS in the hdr file - fixed
Cast 034 - slightly (~0.5nm) off station, another ship anchored on station
Cast 037 - hdr mislabeled cast 036 - fixed
Cast 042 - large swell but no wind
Cast 045 - early prodo, no secchi - estimated at 21m, prodo depths not used
Cast 046 - cast restarted in water
Cast 049 - bottle 16 (43m) tripped as CTD was starting up to 30m
Cast 050 - post-cruise, during intial data processing the secondary oxygen sensor was found to
    be faulty. Do not use CTD secondary oxygen sensor data from casts 051-106. Not even bottle-corrected data.
Cast 051 - 515m bottle may not have fired although tripped, mrk was 0 at 440m so two bottles
    were tripped at 440m. Sequence off, bottle data will sort out depths bottles closed.
Cast 059 & 059A - Santa Barbara Basin; deck unit non-confirmation problem with deep bottles
    CTD data okay but cast redone - 59A is aborted cast with good CTD data, 59 is the 2nd cast
    with good deep bottle confirmations AND data. No cause-effect-solution recorded. Redoing the
    cast seemed enough to fix the problem. DNF was ctd operator.
Cast 069 - mislabeled cast 066, files edited
Cast 078 - 270m upcast wire angle, bottle depth adjusted to get 270m bottle depth
Cast 079 - postcast deck test
Cast 082 - pre-station squall, ~10ft seas
Cast 083 - secondary oxygen sensor giving bad data, no obvious cause; ship rocking & rolling - rough
Cast 084 - rough, lowered CTD at 30-40m/min
Cast 086 - ISUS removed after cast
Cast 087 - 67.1 47.7 bottom shoaling from 268 to 198, CTD terminal depth adjusted; no ISUS - removed
Cast 089 - 1000m casts begin for MBARI-SECRET line 67 starting at 67.55
Cast 091 - no downcast data! "data archiving" started at terminal depth (SMM at the CTD console)
Cast 094 - a second bottle tripped at 800m by mistake, no samples takenfrom bottle 4
Cast 098 - bottle trip confirmation problem at bottle 2 & 3, jumped to 6, surface was bottle 22.
Cast 099 - pump did not come on, brought aboard, checked, redeployed, still no pump, Jen awakened,
    deck unit rebooted, everything working - SMM did not start reattime data acquisition so system was never
    truly on...midnight watch.
Cast 102 - tons of large jellies; O2 at surface was high ~7.5ml/l; sta 63.52
Cast 104 - CTD operator did not start data archiving till the CTD downcast hit ~280m.
Cast 106 - no bottle confirmation on bottle 3 again; bottle 4 said "bottle fired 3, next 5"

JRW 06/08/2015